# Variational Learning of Disentangled Representations

### v1

@@ Abstract
Disentangled representations enable models to separate factors of variation that are shared across experimental conditions from those that are condition-specific. This separation is essential in domains such as biomedical data analysis, where generalization to new treatments, patients, or species depends on isolating stable biological signals from context-dependent effects. While extensions of the variational autoencoder (VAE) framework have been proposed to address this problem, they frequently suffer from leakage between latent representations, limiting their ability to generalize to unseen conditions. Here, we introduce DISCoVeR, a new variational framework that explicitly separates condition-invariant and condition-specific factors. DISCoVeR integrates three key components: (i) a dual-latent architecture that models shared and specific factors separately; (ii) two parallel reconstructions that ensure both representations remain informative; and (iii) a novel max-min objective that encourages clean separation without relying on handcrafted priors, while making only minimal assumptions. Theoretically, we show that this objective maximizes data likelihood while promoting disentanglement, and that it admits a unique equilibrium. Empirically, we demonstrate that DISCoVeR achieves improved disentanglement on synthetic datasets, natural images, and single-cell RNA-seq data. Together, these results establish DISCoVeR as a principled approach for learning disentangled representations in multi-condition settings.
@@ 1 Introduction
Neural network-based models have shown impressive success in learning informative representations of complex data, and are increasingly used in scenarios where each data point $x\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is associated with a condition label $y\in{1,\dots,K}$. In biological applications, for example, the condition may correspond to a treatment, patient, or species. In biological contexts, for instance, the condition might reflect a treatment, a patient, or a species. The learned representations often support downstream tasks such as domain adaptation and transfer learning (e.g., Pan et al. (2010)), enabling models to generalize from observed conditions to new, related ones. This requires learning disentangled data representations that separate components of $x$ shared across all conditions from those specific to a given condition $y$.
@@
In recent years, generative models have emerged as a powerful class of methods for representation learning, thanks to their ability to model complex data distributions and capture underlying latent structure. Prominent examples include Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), Variational Autoencoders (VAEs) (Kingma and Welling, 2014), and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Among these, VAEs and their extensions have become particularly prominent in applications involving transfer learning and domain adaptation (Akrami et al., 2020; Lovrić et al., 2021; Godinez et al., 2022; Zhang et al., 2023), due to their principled probabilistic formulation and their ability to represent uncertainty in learned features.
@@
Several recent VAE-based approaches have been proposed to integrate data from multiple conditions or sources (Xu et al., 2021; Lotfollahi et al., 2019; Boyeau et al., 2022). However, only few explicitly aim to disentangle shared from condition-specific components (Sohn et al., 2015; Klys et al., 2018; Joy et al., 2020). While these methods offer notable improvements in disentanglement, they either depend on carefully crafted priors, difficult to specify in high-dimensional complex settings like single-cell genomics, or fail to achieve a clean separation, limiting their utility in applications that demand robust generalization across conditions.
@@
In this work, we propose a new framework for learning disentangled representations in multi-condition datasets. Our contributions are as follows: (i) We propose a new method based on two distinct reconstruction objectives combined with adversarial learning. Unlike prior approaches, our framework minimizes reliance on restrictive priors or handcrafted components, relying less restrictive assumptions. (ii) We formulate the task of learning disentangled representations as max-min optimization problem, derive a corresponding objective and give theoretical guarantees for the equilibrium of this objective. (iii) Through experiments on synthetic benchmarks and real-world datasets in computer vision and single-cell biology, we show that our approach consistently achieves better disentanglement of shared and condition-specific structure compared to existing methods.
2 DISCoVeR: DISentangling COmmon and Variant Representations
@@
For the task of learning disentangled representations from multi-condition data, we consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ consisting of inputs $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ collected from associated condition labels $y_{i}\in{1,\dots,K}$. We assume that for each class $k$, corresponding to a study or experimental condition, and that the associated subset $\mathcal{D}_{k}\coloneqq\{x_{i}:y_{i}=k\}$ consists of i.i.d. samples drawn from a class-conditional distribution $p(x|y=k)$.
@@ 2.1 Model assumptions
However, note that in this model, $z$ and $w$ are no longer independent if conditioned also on $x$, that is, $z\not\perp w\mid x,y$.
@@ 2.2 Target posterior structure
In our model, the observed data $x$ is generated from two latent variables, $z$ and $w$. The marginal distribution of $z$ is independent of the condition label $y$, making $z$ condition-invariant, whereas $w$ depends on $y$ and is therefore condition-aware. Our goal is to learn probabilistic data representations such that the marginal distributions of $z$ and $w$ preserve the structure $z\perp y$ and $w\not\perp y$, and therefore yield disentangled representations. For this, we approximate the joint posterior $p_{z,w|x,y}$ via variational inference, enforcing that $z$ remains independent of $y$, while $w$ retains dependence on $y$. Our approach follows the empirical Bayes paradigm (Robbins, 1956), and in particular aligns with the principles of $f$-modeling introduced by Efron (2014), where the prior structure is inferred from the data to ensure consistency with observed marginals.
Ideally we would approximate the true posterior $p_{z,w|x,y}$ with a variational distribution $q_{z,w|x,y}$. However, even under simple variational families such as multivariate Gaussians, capturing dependencies between $z$ and $w$ requires modeling a full covariance structure, which becomes computationally intractable in high dimensions. To mitigate this, we approximate the posterior $p_{z,w|x,y}$ with a factorized variational distribution $q_{z|x}\ q_{w|x,y}$.
Our variational approximation is guided by two complementary objectives: (i) $q_{z|x}$ closely approximating the marginal posterior $p_{z|x}$; and (ii) the product $q_{z|x}\ q_{w|x,y}$ closely approximating the true posterior $p_{z,w|x,y}$. Formally, given variational families [^1] Here we consider general families and specify our concrete choices in Section 2.4. $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ we seek to find $q^{*}_{z|x}\in\mathcal{Q}_{z}$ and $q^{*}_{w|x,y}\in\mathcal{Q}_{w}$ that minimize the following sum of Kullback-Leibler (KL) divergences:
$q^{*}_{z|x},\ q^{*}_{w|x,y}=\arg\min_{\begin{subarray}{c}q_{z|x}\in\mathcal{Q}_{z}\\ q_{w|x,y}\in\mathcal{Q}_{w}\end{subarray}}\operatorname{D_{\mathrm{KL}}}(q_{z|x}\ \|\ p_{z|x})+\operatorname{D_{\mathrm{KL}}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y}).$ (2)
@@ 2.3 Optimization objective
Direct computation of the KL divergences is Equation 2 is intractable. Therefore, we introduce a surrogate objective composed of three terms, two of which are evidence lower bound (ELBO) terms that minimize the sum of KL divergences without requiring their direct evaluation, and an additional adversarial term.
The corresponding ELBO objective to minimize $\operatorname{D_{\mathrm{KL}}}(q_{z|x}\ \|\ p_{z|x})$ is
$\mathcal{L}_{z}(q_{z|x},p;x)\coloneqq\mathbb{E}_{q_{z|x}}\left[\log p(x\mid z)\right]-\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z}),$ (3)
and the ELBO objective for the second KL term, $\operatorname{D_{\mathrm{KL}}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y})$, is given by
$\textstyle\mathcal{L}_{w}(q_{w|x,y},p;x,y)\!\coloneqq$ $\mathbb{E}_{q_{z|x}}\!\left[\mathbb{E}_{q_{w|x,y}}\!\left[\log\left(p(x|z,w)\right)\right]\right]\!\!-\!\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z}\right)\!\!-\!\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (4)
Note that $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ is the ELBO objective that corresponds to a factorized posterior $q_{z|x}q_{w|x,y}$. In Proposition 2.1 we examine the gap between this objective, and an ELBO term corresponding to a full variational posterior $q_{z,w|x,y}$. This can be interpreted as the cost of enforcing a condition-invariant latent representation, specifically, constraining $z$ to depend only on $x$.
@@
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right).$
where $\operatorname{ELBO}(q,p;x,y)\coloneqq\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$.
@@
Note that a full definition of the objectives in Equations 3 and 4 requires the specification of corresponding prior distributions, namely $p_{z}$ and $p_{w|y}$. We defer their definitions to Section 2.4.
The ELBO objective in Equation 4 is defined over the conditional log-likelihood $\log p(x\mid y)$, and requires the addition of $\log p(y)$ to correspond to the log-marginal likelihood $\log p(x,y)$. However, since $\log p(y)$ does not depend on the variational distributions, including it in the objective has no effect on optimization. Still, we wish to ensure that the implicit marginal over $y$—as inferred from the latent representations, matches the true marginal $p(y)$. To this end, we introduce an auxiliary classifier $g(y\mid z)$ as a form of posterior regularization (Ganchev et al., 2010). This classifier captures the residual predictive signal about $y$ contained in $z$, and is trained via the expected negative log-likelihood $-\mathbb{E}_{q(z\mid x)}\log g(y\mid z)$. If $z$ is indeed independent of $y$, then $g(y\mid z)$ will approximate the marginal distribution $p(y)$. By penalizing deviations from this behavior, we enforce the structural constraint $z\perp y$ in the learned representation.
Note that for this term to effectively encourage $q_{z|x}$ to discard condition-specific information, the classifier $g_{y|z}\in\mathcal{G}$ must be optimized separately from the the rest of the model. This prevents degenerate solutions in which the third term is minimized without truly removing information about $y$ from $z$, for example, by collapsing $g$ to always predict a constant output, regardless of $z$.
@@
$-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]-2\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (6)
Finally, to enable flexible trade-offs between reconstruction expressiveness and disentanglement, we introduce weighting terms $\alpha=(\alpha_{1},\alpha_{2})$ into the objective following the motivation of $\beta$-VAEs (Higgins et al., 2017):
@@
$-\mathbb{E}_{q_{z|x}}\log g(y\mid z)-\alpha_{1}\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})-\alpha_{2}\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$
@@ 2.4 Latent prior models and variational approximations
For the condition-aware latent variable $w$, we define a class-conditional Gaussian prior $p_{w|y}$. As a simple specification, we model $z$ and $w$ to be of the same dimension, and specify
$p(w\mid y=k)=\mathcal{N}(\mu_{k},I)$ (9)
$\mu_{k}\coloneqq\mathbb{E}_{p_{x|y=k}}\left[\mathbb{E}_{q_{z|x}}[z]\right].$ (10)
Here $\mu_{k}$ is the mean of the inferred latent representations $z$ within each class. Similarly, for different dimensions of $z$ and $w$ the mean aggregation can be replaced with a neural network that maps the inferred representations $z$ for each class to the parameters of the Gaussian prior.
This specification induces a coupling between the two latent variables through the data distribution. Aligning $p_{w|y}$ with the class-wise expectations of the invariant variable, further encourages $q_{z|x}$ to encode informative representations, since capturing the shared structures will now not only increase $\mathcal{L}_{z}(q_{z|x},p;x)$, but also decrease $\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x,y}\,\|\,p_{w|y}\right)$, and thus increase $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$.
@@ 3 Encoder-decoder model
In order to optimize the objective in Equation 8 with respect to $q_{z|x},q_{w|x,y}$ and $g_{y|z}$ over the dataset $\mathcal{D}$, we introduce an encoder-decoder framework (for illustration see Figure 1(b)) in which two separate reconstructions of $x$ are generated: one, denoted $\hat{x}\sim p_{x|z}$, where $z$ is sampled from the condition-invariant posterior $q_{z|x}$, and the other, denoted $\tilde{x}\sim p_{x|z,w}$, where in addition, $w$ is sampled from the condition-aware posterior $q_{w|x,y}$. The corresponding algorithm is summarized in Algorithm 1.
@@
Adversarial classifier   To avoid constructing a complex classifier directly from $z$ to $y$, we leverage the reconstruction $\hat{x}$ from $z$. As a proxy for $g(y|z)$, we instead predict $y$ from $\hat{x}$, allowing the use of a simple logistic regression on $\hat{x}$. Therefore, we define the adversarial classifier as a multinomial logistic regression $g_{\beta}:\hat{\mathcal{X}}\rightarrow[0,1]^{K}$, parametrized by class-specific weight vectors $\beta=\{\beta_{k}\}_{k=1}^{K}$.
While, from an information-theoretic perspective, this substitution weakens the estimation of the cross-entropy term $-\mathbb{E}_{q_{z|x}}\log g(y\mid z)$, since by the data processing inequality, $\hat{x}$ can contain at most the information present in $z$, we observed this to be often preferable from an optimization perspective. Predicting $y$ from $\hat{x}$ helps reduce the variance introduced by sampling $z\sim q_{\theta_{z}}$, and thus acts as a form of regularization that prevents $q_{\theta_{z}}$ from overfitting to noisy penalties produced by the classifier.
Condition aware representation   A labeled input pair $(x,y)\in\mathcal{X}\times\{1,\dots,K\}$ is mapped to the parameters $\theta_{w}=(\mu_{w},\sigma^{2}_{w})$ using an encoder neural network neural-network $f^{w}_{\rho}:\mathcal{X}\times\{1,\dots,K\}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\rho$. A sample $w\sim q_{\theta_{w}}$ is then drawn, and the pair $(z,w)$ is mapped to a reconstruction $\tilde{x}$ via a decoder neural-network $h^{z,w}_{\eta}:\mathbb{R}^{d}\rightarrow\mathcal{X}$, parametrized by weights $\eta$. To compute the prior $p_{w|y}$, we estimate the class-specific mean as $\hat{\mu}_{k}\coloneqq\frac{1}{n_{k}}\sum_{i;y_{i}=k}z_{i}$ where each $z_{i}\sim p(z\mid x_{i})$ is sampled from the encoder given an input $x_{i}$ with label $y_{i}=k$, and $n_{k}$ is the number of training points with the label $y=k$.
@@
In practice, following standard approaches in VAE-based models, we make two approximations: (i) we use a single-sample Monte Carlo estimate to approximate the expectations in Equation 8, and (ii) instead of directly sampling from $q_{\theta}$, we employ the reparameterization trick to enable differentiable sampling. Specifically, we sample $\epsilon\sim\mathcal{N}(0,I)$ and obtain a sample from $q_{\theta}$ by applying a deterministic transformation of $\epsilon$ based on the variational parameters $\theta$.

### v2

@@ Abstract
Disentangled representations allow models to separate factors shared across conditions from those that are condition-specific. This separation is crucial in domains such as biomedicine, where generalization to new treatments, patients, or species requires isolating stable biological signals from context-dependent effects. While several VAE-based extensions aim to achieve this, they often exhibit leakage between latent variables, limiting generalization. We introduce DisCoVR, a variational framework that explicitly separates invariant and condition-specific factors through: (i) a dual-latent architecture, (ii) parallel reconstructions to keep both representations informative, and (iii) a max–min objective that enforces separation without handcrafted priors. We show this objective maximizes data likelihood, promotes disentanglement, and admits a unique equilibrium. Empirically, DisCoVR achieves stronger disentanglement on synthetic data, natural images, and single-cell RNA-seq datasets, establishing it as a principled approach for multi-condition representation learning.
@@ 1 Introduction
Neural network–based models excel at learning rich representations of complex data and are increasingly applied in settings where each data point $x\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is associated with a condition label $y\in{1,\dots,K}$. In biology, for example, conditions may represent treatments, patients, or species. Such representations are valuable for tasks like domain adaptation and transfer learning (transfer), where models must generalize from observed to novel conditions. Achieving this requires disentangled representations that separate factors shared across conditions from those specific to each $y$.
@@
Generative models provide a natural framework for uncovering latent structure, with prominent examples including Generative Adversarial Networks (GANs) (goodfellow2020generative), Variational Autoencoders (VAEs) (vae), and diffusion models (sohl2015deep; ho2020denoising). Among these, VAEs and their extensions are particularly well-suited to transfer learning and domain adaptation (akrami2020brain; lovric2021should; godinez2022design; zhang2023variational), thanks to their probabilistic foundation and ability to capture uncertainty.
@@
Thus, several VAE-based methods have been proposed to integrate data across multiple conditions or sources (xu2021probabilistic; lotfollahi2019scgen; boyeau2022deep), but only a few explicitly disentangle shared and condition-specific components (CVAE; CSVAE; CCVAE). While these approaches improve separation to some degree, they often depend on handcrafted priors that are difficult to design in high-dimensional domains like single-cell genomics, and they frequently exhibit residual leakage between latent varaibles, limiting generalization across conditions.
@@
In this work, we introduce a framework for learning disentangled representations in multi-condition datasets. Our main contributions are: (i) a method combining two distinct reconstruction objectives with adversarial learning, reducing reliance on restrictive priors or handcrafted components; (ii) a max–min formulation of disentangled representation learning, along with a corresponding objective and theoretical guarantees for its equilibrium; and (iii) through experiments on synthetic benchmarks and real-world datasets, we show that our approach consistently improves upon existing methods disentanglement of shared and condition-specific structure.
2 DisCoVR: Disentangling Common and Variant Representations
@@
For the task of learning disentangled representations from multi-condition data, we consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ consisting of inputs $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ collected from associated condition labels $y_{i}\in\{1,\dots,K\}$. For each class $k$ (corresponding to a study or experimental condition), the associated subset $\mathcal{D}_{k}\coloneqq\{x_{i}:y_{i}=k\}$ consists of i.i.d. samples drawn from a class-conditional distribution $p(x\mid y=k)$.
@@ 2.1 Model assumptions
However, note that in this formulation, $z$ and $w$ are no longer independent if conditioned also on $x$, that is, $z\not\perp w\mid x,y$.
@@ 2.2 Target posterior structure
In our model, each observation $x$ is generated from two latent variables: $z$, which is condition-invariant, and $w$, which is condition-aware through its dependence on $y$. Our goal is to learn probabilistic representations where the marginals of $z$ and $w$ preserve this structure, yielding disentangled factors. Thus, we seek to approximate the posterior $p_{z,w\mid x,y}$.
However, approximating the full posterior with a variational distribution $q_{z,w|x,y}$ is intractable: even for simple variational families such as Gaussians, modeling the dependencies between $z$ and $w$ requires a full covariance structure, which is computationally prohibitive in high dimensions. To mitigate this, we employ a factorized approximation $q_{z|x}\ q_{w|x,y}$.
Our variational approximation is guided by two complementary objectives: (i) $q_{z|x}$ closely approximating the marginal posterior $p_{z|x}$; and (ii) the product $q_{z|x}\ q_{w|x,y}$ closely approximating the true posterior $p_{z,w|x,y}$. Formally, given variational families [^1] Here we consider general families and specify our concrete choices in §2.4. $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ we seek to find $q^{*}_{z|x}\in\mathcal{Q}_{z}$ and $q^{*}_{w|x,y}\in\mathcal{Q}_{w}$ that minimize the following sum of Kullback-Leibler (KL) divergences:
$q^{*}_{z|x},\ q^{*}_{w|x,y}=\arg\min_{\begin{subarray}{c}q_{z|x}\in\mathcal{Q}_{z}\\ q_{w|x,y}\in\mathcal{Q}_{w}\end{subarray}}D_{\mathrm{KL}}(q_{z|x}\ \|\ p_{z|x})+D_{\mathrm{KL}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y}).$ (2)
@@ 2.3 Optimization objective
Since direct evaluation of the KL divergences in Equation 2 is intractable, we optimize a surrogate objective consisting of two ELBO terms.
The corresponding ELBO objective to minimize $D_{\mathrm{KL}}(q_{z|x}\ \|\ p_{z|x})$ is
$\mathcal{L}_{z}(q_{z|x},p;x)\coloneqq\mathbb{E}_{q_{z|x}}\left[\log p(x\mid z)\right]-D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z}),$ (3)
and the ELBO objective for the second KL term, $D_{\mathrm{KL}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y})$, is given by
$\textstyle\mathcal{L}_{w}(q_{w|x,y},p;x,y)\!\coloneqq$ $\mathbb{E}_{q_{z|x}}\!\left[\mathbb{E}_{q_{w|x,y}}\!\left[\log\left(p(x|z,w)\right)\right]\right]\!\!-\!D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z}\right)\!\!-\!D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (4)
Note that $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ is the ELBO objective that corresponds to a factorized posterior $q_{z|x}q_{w|x,y}$. In Proposition 2.1 we examine the gap between this objective, and an ELBO term corresponding to a full variational posterior $q_{z,w|x,y}$. This can be interpreted as the cost of enforcing a condition-invariant latent representation, specifically, constraining $z$ to depend only on $x$. Proposition
@@
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)=\mathbb{E}_{q_{w|x,y}}\left[\text{KL}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right)\right].$
where $\operatorname{ELBO}(q,p;x,y)\coloneqq\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$.
@@
Note that a full definition of the objectives in Equations 3 and 4 requires the specification of corresponding prior distributions, namely $p_{z}$ and $p_{w|y}$. We defer their definitions to §2.4.
Equation 4 provides an evidence lower bound on the conditional log-likelihood $\log p(x\mid y)$. By adding $\log p(y)$, this bound extends to the joint log-marginal likelihood $\log p(x,y)$. Beyond optimizing this objective, we aim to ensure that the marginal distribution over $y$ implicitly induced by the latent representations is consistent with the true $p(y)$.
To this end, we introduce an auxiliary classifier $g(y\mid z)$ as a form of posterior regularization (ganchev2010posterior). This classifier captures the residual predictive signal about $y$ in $z$ and is trained by minimizing the expected negative log-likelihood $\mathbb{E}_{q(z\mid x)}\log g(y\mid z)$. If $z$ is truly independent of $y$, then $g(y\mid z)$ will approximate the marginal distribution $p(y)$. By penalizing deviations from this behavior, we enforce the structural constraint $z\perp y$ in the learned representation.
Note that for this term to effectively encourage $q_{z|x}$ to discard condition-specific information, the classifier $g_{y|z}\in\mathcal{G}$ must be trained adversarially, with its own update step. This prevents degenerate solutions in which the loss is minimized without removing information about $y$ from $z$, for example, by collapsing $g$ to a constant predictor that ignores its input.
@@
$-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]-2D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (6)
Finally, to enable flexible trade-offs between reconstruction expressiveness and disentanglement, we introduce weighting terms $\alpha=(\alpha_{1},\alpha_{2})$ into the objective following the motivation of $\beta$-VAEs (beta_vae):
@@
$-\mathbb{E}_{q_{z|x}}\log g(y\mid z)-\alpha_{1}D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})-\alpha_{2}D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$
@@ 2.4 Latent prior models and variational approximations
For the condition-aware latent variable $w$, we define a class-conditional Gaussian prior $p_{w|y}$. As a simple choice, we let $w$ have the same dimensionality as $z$ and specify
$p(w\mid y=k)=\mathcal{N}(\mu_{k},I),\qquad\mu_{k}\coloneqq\mathbb{E}_{p_{x|y=k}}\left[\mathbb{E}_{q_{z|x}}[z]\right].$ (9)
Here $\mu_{k}$ is the mean of the inferred latent representations $z$ within the $k$-th class [^2] Similarly, for different dimensions of $z$ and $w$ the mean aggregation can be replaced with a neural network that maps the inferred representations $z$ for each class to the parameters of the Gaussian prior..
This specification induces a coupling between the two latent variables through the data distribution. Aligning $p_{w|y}$ with the class-wise expectations of the invariant variable, further encourages $q_{z|x}$ to encode informative representations, since capturing the shared structures will now not only increase $\mathcal{L}_{z}(q_{z|x},p;x)$, but also decrease $D_{\mathrm{KL}}\left(q_{w\mid x,y}\,\|\,p_{w|y}\right)$, and thus increase $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$.
@@ 3 Encoder-decoder model
In order to optimize the objective in Equation 8 with respect to $q_{z|x},\,q_{w|x,y}$, and $g_{y|z}$ over the dataset $\mathcal{D}$, we introduce an encoder-decoder framework (illustrated in Figure 1(b)). In this framework, two separate reconstructions of $x$ are generated: one, denoted $\hat{x}\sim p_{x|z}$, where $z$ is sampled from the condition-invariant posterior $q_{z|x}$, and the other, denoted $\tilde{x}\sim p_{x|z,w}$, where in addition, $w$ is sampled from the condition-aware posterior $q_{w|x,y}$. The corresponding algorithm is summarized in Algorithm 1.
@@
Adversarial classifier   Rather than learning a complex classifier from $z$ to $y$, we use the reconstruction $\hat{x}$ from $z$, and predict $y$ from $\hat{x}$ via a simple multinomial logistic regression $g_{\beta}:\hat{\mathcal{X}}\rightarrow[0,1]^{K}$, with class-specific weights $\beta={\beta_{k}}_{k=1}^{K}$, or a shallow MLP. Since $\hat{x}$ is a deterministic function of $z$, this is equivalent to applying a restricted classifier on $z$. By the data processing inequality, such a classifier can only capture a subset of the information $z$ contains about $y$; thus, maximizing this lower bound on $I(z;y)$ also maximizes $I(z;y)$ itself. Although this substitution weakens the estimation of the cross-entropy term $\mathbb{E}{q{z\mid x}}\log g(y\mid z)$, from an information-theoretic standpoint, we observed this to be often advantageous in practice: predicting $y$ from $\hat{x}$ reduces the variance introduced by sampling $z\sim q_{\theta_{z}}$, providing a regularizing effect that prevents $q_{\theta_{z}}$ from overfitting to noisy classifier signals.
Condition aware representation   A labeled input pair $(x,y)\in\mathcal{X}\times\{1,\dots,K\}$ is mapped to the parameters $\theta_{w}=(\mu_{w},\sigma^{2}_{w})$ using an encoder neural-network $f^{w}_{\rho}:\mathcal{X}\times\{1,\dots,K\}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\rho$. A sample $w\sim q_{\theta_{w}}$ is then drawn, and the pair $(z,w)$ is mapped to a reconstruction $\tilde{x}$ via a decoder neural-network $h^{z,w}_{\eta}:\mathbb{R}^{d}\rightarrow\mathcal{X}$, parametrized by weights $\eta$. To compute the prior $p_{w|y}$, we estimate the class-specific mean as $\hat{\mu}_{k}\coloneqq\frac{1}{n_{k}}\sum_{i;y_{i}=k}z_{i}$ where each $z_{i}\sim p(z\mid x_{i})$ is sampled from the encoder given an input $x_{i}$ with label $y_{i}=k$, and $n_{k}$ is the number of training points with the label $y=k$.
@@
In practice, following standard approaches in VAE-based models, (i) we use a single-sample Monte Carlo estimate to approximate the expectations in Equation 8, and (ii) instead of directly sampling from $q_{\theta}$, we employ the reparameterization trick to enable differentiable sampling. Specifically, we sample $\epsilon\sim\mathcal{N}(0,I)$ and obtain a sample from $q_{\theta}$ by applying a deterministic transformation of $\epsilon$ based on the variational parameters $\theta$.